Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 543
Entry time: 28 Aug 2024, 09:52
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 28.08.2024 |
|
|
General Announcement:
- Communications Group thanks all people involved on the SRI visit. The extension of the beam delivery also worked nicely to instruments.
- As PRC, A. Galler says so far there is nothing to report or discuss for the current week.
- For next Friday the UOM will have presentations from the instruments (on stability) and from the FD activities.
CTRL:
- Nothing to report.
DA:
- All fine with DA.
EEE:
- Not present.
VAC:
- Not present.
XPD:
- No issues; preparing for the FD studies;
- ZZ in XTD2 on Friday to repair camera needed for XFLO.
DET:
- MID AGIPD HV issue on weekend: no power glitch happened that was registered from Desy systems;
still not understood the reason but likely a local power glitch (small voltage drop (peak) was seen) and will be further tracked.
ITDM:
- E-mail sent yesterday for general configuration on displays.
XRO:
- Not present.
LAS:
- SA1-PPL Amphos service ongoing;
- SA3-PPL changing repetition rate for next experiments.
UNSYS:
- Everything is fine.
SPB:
- Nothing to report.
- Issues with trigger distribution systems yesterday happened but was fixed on the fly.
FXE:
- Nothing to report.
MID:
- Nothing to report.
HED:
- Nothing to report.
SCS:
- Not present.
SQS:
- Not present.
SXP:
- Not present. |